40m
QIL
Cryo_Lab
CTN
SUS_Lab
TCS_Lab
OMC_Lab
CRIME_Lab
FEA
ENG_Labs
OptContFac
Mariner
WBEEShop
|
40m Log |
Not logged in |
 |
|
Thu Nov 7 14:39:49 2013, Jenne, Update, LSC, Lock Acquisition Game Plan
|
Fri Nov 8 18:12:21 2013, Jenne, Update, LSC, PRFPMI: Not crossing any resonances 
|
Tue Nov 12 16:49:22 2013, Jenne, Update, LSC, Xend QPD and Whitening board pulled   
|
Wed Nov 20 17:05:15 2013, Jenne, Update, LSC, Xend QPD and Whitening board replaced
|
Wed Dec 18 00:26:15 2013, Jenne, Update, LSC, Xend QPD schematic investigation
|
Wed Dec 18 11:32:34 2013, rana, Update, LSC, Xend QPD schematic investigation
|
Wed Nov 13 01:35:40 2013, Jenne, Update, LSC, PRM motion causing trouble? 
|
Wed Nov 13 09:31:15 2013, Gabriele, Update, LSC, PRM motion causing trouble?
|
Wed Nov 13 19:22:58 2013, Jenne, Update, LSC, PRM motion correlated to intracavity power 
|
Thu Nov 14 02:50:43 2013, Jenne, Update, LSC, PRM oplev measured and modeled TF  
|
Mon Nov 18 21:02:54 2013, Jenne, Update, LSC, PRM oplev measured and modeled TF 
|
Mon Nov 18 21:40:24 2013, Koji, Update, LSC, PRM oplev measured and modeled TF
|
Tue Nov 19 14:47:44 2013, Jenne, Update, LSC, PRM oplev measured and modeled TF
|
Thu Nov 14 02:55:26 2013, rana, Update, SUS, PRM motion correlated to intracavity power
|
Thu Nov 14 11:41:19 2013, Steve, Update, SUS, PRM sensors effected by IR
|
Thu Nov 14 14:35:12 2013, Steve, Update, SUS, IR effect on MC and PRM sensors
|
Fri Dec 6 10:03:07 2013, Steve, Update, SUS, IR effect on MC sensors only 
|
|
Message ID: 9384
Entry time: Thu Nov 14 11:41:19 2013
In reply to: 9383
Reply to this: 9386
|
Author: |
Steve |
Type: |
Update |
Category: |
SUS |
Subject: |
PRM sensors effected by IR |
|
|
IR off for 11 minutes. The PRM face sensors are effected. The PRM side and the rest of the SUS OSEMS are not effected.
|
|